Find the equation of the circle, in general form, that has endpoints of a diameter at the points (5, 3) and (0, 2).
Answer by stanbon(75887)

Find the equation of the circle, in general form, that has endpoints of a diameter at the points (5, 3) and (0, 2).
The center is half-way between so is at (5/2 , 1)
The radius is the distance from the center to one of the end-points.
radius = sqrt((5/2)-0)^2 + (1-2)^2)= sqrt((25/4)+1)=sqrt(29/4)
Equation form is (x-(5/2))^2 + (y-1)^2= 29/4
Cheers,
Stan H.